OpenStudy (aravindg):

How do we get average energy density as \[\LARGE \dfrac{1}{2} \epsilon_0 E^2\]

OpenStudy (aravindg):

\[\LARGE \dfrac{1}{2} \epsilon_0 E^2\]

OpenStudy (aravindg):

@Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

The easiest way is to rewrite electric energy in a capacitor EE = (1/2)CU² using electric field E = U/d and capacitance C = \(\epsilon_0\)A/d You will end up with EE = \(\dfrac {\epsilon_o\,E^2}{2}\times Ad\) giving density \(u_e=\dfrac {\epsilon_o\,E^2}{2}\) Full derivation can be obtained from conservation of energy carried by an electromagnetic wave from Maxwell's equations, but requires more complicated differential mathematics.